# Positive operator-valued noncommutative polynomials are squares

### v1

@@ Abstract.
The key technical input is a canonical tuple of self-adjoint operators $A$ and unitaries $U$, respectively, constructed from the left-regular representation on Fock space. We prove that, up to a universal constant, the norms $\|p(A)\|$ and $\|p(U)\|$ bound the operator norm of any positive semidefinite Gram matrix $S$ representing the sos polynomial $p$. As a consequence, the cone of (sums of) squares of polynomials is closed in the weak operator topology. Exploiting this closedness, the GNS construction associates to a separating linear functional a finite-rank positive semidefinite noncommutative Hankel matrix and, on its range, produces the desired tuple $Y$.
@@ 1.1. Notation
Equip $\mathcal{A}$ with the involution as follows. On letters, $x_{j}^{*}=x_{j},$ on a word $w=x_{i_{1}}\cdots x_{i_{n}}\in\langle x\rangle,$
@@ 1.1.1. Trigonometric polynomials
for all $p\in\mathscr{A}$. The notions of length of a word, degree of a polynomial, etc. extend naturally to $\mathscr{A}$ and we let
@@ 1.2. Main results
There are several novel results and constructions of independent interest necessitated by the consideration of operator coefficients. In the large, the proofs of Theorem 1.1 and Theorem 1.3 proceed in the by a now standard approach to establishing sum of squares (sos) representations (factorizations). Namely, Hahn-Banach separation produces a separating linear functional $\varphi$ followed by a Gelfand-Naimark-Segal (GNS) construction based on $\varphi$ that ultimately produces a tuple $Y$. Here we roughly follow the outline of [MP05]. A key construction is that of a tuple of self-adjoint operators $A$ based upon the left regular representation on Fock space. See Section 3. It is then shown, that, up to a universal constant, for a sums of squares polynomial $p,$ up to a universal constant, the norm of $p(A)$ bounds the norm of any non-commutative psd Gram matrix $S$ that represents $p.$ See Proposition 3.3. Via this route we obtain the key fact needed for Hahn-Banach separation. Namely that the cone of sums of squares is weak operator topology (WOT) closed. In the GNS construction, we introduce a new argument that exploits the WOT to associate to a separating linear functional a finite rank psd non-commutative representing Hankel matrix. Continuing with the GNS construction, a proof of Theorem 1.3 demands a new construction to produce a tuple of unitary operators together with a representing vector that realizes the separating linear functional as a vector state. See Subsection 7.2.∎
@@ 1.3. Reader’s guide
The paper is structured as follows. The convex cone of sums of squares (making an appearance in (1.3) of Theorem 1.1) is introduced and characterized in the next Section 2. In Section 3 we define creation operators $L_{i}$ on the full Fock $\mathcal{F}_{\tt{g}}^{2}$ space and their symmetrized analogs $A_{i}$. How they pertain to the sum of squares statement at hand is explored in Subsection 3.3, where evaluations at $A$ are used to extract coefficients of a polynomial. In Section 4 we introduce a suitable topology on $\mathcal{A}_{d}$ and collect all the necessary topological properties needed in the sequel. With respect to this topology, the convex cone of sums of squares is closed, see Proposition 4.4. The fact that the cone is closed allows for an application of the Hahn–Banach Separation Theorem, which is then followed by an appropriate version of the GNS construction, carried out in Section 5 See Proposition 5.3. Then Theorem 1.1 is proved in Section 6 and Theorem 1.3 is proved in Section 7. The paper concludes with an Appendix that collects auxiliary results that may be well-known to experts.
@@ 2. Convex Cone of (Sums of) Squares
In this section, a key player in the proof of Theorem 1.1, the convex cone of sums of squares of polynomials, is introduced and studied. The main result in this section is Proposition 2.2, which gives a bound on the number of sums of squares needed to write a polynomial as a sum of squares.
@@
so that $r^{*}r$ has a representation as $V_{d}^{*}SV_{d}$ for a psd matrix $S.$
@@
A polynomial $p\in\mathcal{A}_{2d}$ is in $\mathcal{C}_{d}$ if and only if there is a psd block matrix $S$ such that
$p\ =\ V_{d}^{*}SV_{d}.$ (2.2)
In fact, if $p=V_{d}^{*}SV_{d},$ then factoring $S=\sum_{j=1}^{N(d)}R_{j}R_{j}^{*}$ with $R_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ as in Lemma 2.1, setting $r_{j}=R_{j}^{*}V_{d}$ gives,
@@
We call any psd block matrix satisfying equation (2.2) a Gram representation for $p.$
@@
Given a sum of squares $p=\sum_{i=1}^{N(d)}r_{i}^{*}r_{i},$ writing $r_{j}=R_{j}^{*}V_{d}$ gives $p=\sum_{i=1}^{N(d)}V_{d}^{*}R_{i}R_{i}^{*}V_{d},$ where $R_{i}$ is the coefficient vector corresponding to the polynomial $r_{i}.$ It follows that $p=V_{d}^{*}SV_{d},$ where $S=\sum_{i=1}^{N(d)}R_{i}R_{i}^{*}.$ In particular, $S:\mathcal{H}^{N(d)}\to\mathcal{H}^{N(d)}$ is a psd linear map.
Conversely, suppose there is a psd linear map $S:\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that $p=V_{d}^{*}SV_{d}.$ By Lemma 2.1, there exist $R_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that $S=\sum_{j=1}^{N(d)}R_{j}R_{j}^{*}.$ Setting $r_{j}=R_{j}^{*}V_{d}$, one obtains $p=\sum_{j=1}^{N(d)}r_{j}^{*}r_{j}.$
By what has already been proved, if $p,q\in\mathcal{C}_{d},$ then there exist (psd) Gram representations $p=V^{*}S_{p}V$ and $q=V^{*}S_{q}V.$ Now $p+q=V_{d}^{*}(S_{p}+S_{q})V_{d}.$ Since $S_{p}+S_{q}:\mathcal{H}^{N(d)}\to\mathcal{H}^{N(d)}$ is a psd linear map, what has already been proved shows $p+q\in\mathcal{C}_{d}.$
@@
$\mathcal{C}_{d}=\left\{V_{d}^{*}SV_{d}:\quad S=[S_{v,w}]_{v,w\in\langle x\rangle_{d}}\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)},\quad S\succeq 0\right\}.$
@@ 3. Full Fock Space and Gram Matrices
This section recalls the well-known definition of the full Fock space [AP95, JMS21], the creation operators [Fra84], and introduces their symmetrized variants in (3.2). In Subsection 3.3 we explore how these self-adjoint operators are used to extract the coefficients of a polynomial. The main result is Proposition 3.3 showing that the set of positive semidefinite Gram matrices of a polynomial is norm bounded.
@@ 3.2. Left creation operators
Clearly, each $L_{i}$ is an isometry. Moreover $L_{i}^{*}L_{j}=0$ if $i\neq j.$ Thus the tuple $L=(L_{1},\dots,L_{{\tt{g}}})$ is a row isometry. Let
$A_{i}\ =\ L_{i}+L_{i}^{*},\quad i=1,\dots,{\tt{g}}.$ (3.2)
@@
The proof proceeds by induction on $|w|.$ If $w$ is the empty word, then $A^{w}\Omega=\Omega=e_{\varnothing}.$ Assume the claim holds for all words of length $\leq n-1$. Let $w=x_{i_{1}}\dots x_{i_{n}}$ be of length $n.$ Then
$A^{w}\Omega\ =$ $\ (L_{i_{1}}+L_{i_{1}}^{*})A_{i_{2}}\cdots A_{i_{n}}\Omega$
$\ =$ $\ L_{i_{1}}A_{i_{2}}\cdots A_{i_{n}}\Omega+L_{i_{1}}^{*}A_{i_{2}}\cdots A_{i_{n}}\Omega$
$\ =$ $\ L_{i_{1}}A^{\tilde{w}}\Omega+L_{i_{1}}^{*}A^{\tilde{w}}\Omega,$
where $\tilde{w}=x_{i_{2}}\dots x_{i_{n}}$ is a word of length n1.n-1. By the induction hypothesis,
@@
The $N(d)\times N(d)$ scalar matrix
$M_{d}\ =\ \begin{bmatrix}\langle A^{w}\Omega,e_{v}\rangle\end{bmatrix}_{v,w\in\langle x\rangle_{d}}$
@@
Recall that we have endowed $\langle x\rangle$ with graded lexicographic order. If $v>w$, then
@@
by Lemma 3.1. Hence $M_{d}$ is an upper triangular matrix. Moreover, each diagonal entry is 11 by Lemma 3.1. Thus, $M_{d}$ is invertible. ∎
@@ 3.3. Extraction formula for coefficients
Let $q=\sum Q_{w}w\in\mathcal{A}_{d}.$ For any basis vector $e_{v}\in\mathcal{\mathcal{F}}^{2}_{{\tt{g}}}$, define the linear functional
$\Omega_{v}:\mathcal{B}(\mathcal{F}^{2}_{{\tt{g}}})\to\mathbb{C},\qquad\Omega_{{v}}(T)=\langle T\Omega,e_{v}\rangle.$
@@
$\ =$ $\ \sum\limits_{w}\langle A^{w}\Omega,e_{v}\rangle\,Q_{w}\ =\ \sum\limits_{w}[M_{d}]_{v,w}Q_{w},$
where $M_{d}]_{v,w}$ is the $(v,w)$ entry of the matrix $M_{d}.$ In short,
$Z(q)\ =\ M_{d}Q,$ (3.3)
where $Z(q)$ and $Q$ are column vectors with $Z_{v}(q)$ and $Q_{v}$ as the $v^{\rm th}$ entry of $Z$ and $Q,$ respectively. Since, by Lemma 3.2, $M_{d}$ is invertible,
$Q\ =\ M_{d}^{-1}Z(q).$ (3.4)
We refer to equation (3.4) as the extraction formula for the coefficients of $q.$ Note that this formula depends only upon $q(A);$ that is, the coefficients of $q$ are determined uniquely by $q(A).$
It follows from equation (3.4) that there exists a positive constant $\lambda_{d}$ (independent of $q$) such that
$\|Q_{w}\|\ \leq\ \lambda_{d}\,\|q(A)\|\quad\text{ for all $w\in\langle x\rangle_{d}$.}$ (3.5)
@@
$\Gamma_{p}\ =\ \{\,S\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}\ :\quad S\succeq 0,\quad V_{d}^{*}SV_{d}=p\,\}$
is norm bounded ((with respect to the operator norm on $\mathcal{B}(\mathcal{H}^{N(d)})$)). More precisely, there exists a constant $\mu_{d}$ ((depending only on $d$ and $\tt{g}$} and not on $p$)) such that, for all $S\in\Gamma_{p},$
$\|S\|\ \leq\ \mu_{d}\;\|p(A)\|,$
where the tuple $A$ is defined in (3.2).
@@
Fix $p\in\mathcal{C}_{d}$ and $S\in\Gamma_{p}.$ Thus $p=V_{d}^{*}SV_{d}.$ By Proposition 2.2, there exists $Q_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that
@@
$\sum_{u,v\in\langle x\rangle_{d}}\|S_{u,v}\|\ \leq\ \sum_{u,v\in\langle x\rangle_{d}}\,\sum_{j=1}^{N(d)}\|Q_{j,u}Q_{j,v}^{*}\|\ \leq\ N(d)^{3}\lambda_{d}^{2}\;\|p(A)\|.$
It follows that $\|S\|\leq\mu_{d}\,\|p(A)\|$ for $\mu_{d}=N(d)^{3}\lambda_{d}^{2}.$
4. Topology on $\mathcal{A}_{d}$
The main purpose of this section is to define a well-behaved topology on $\mathcal{A}_{2d}$ in which the convex cone of sums of squares $\mathcal{C}_{d}$ is closed. The main property of this topology is given in Proposition 4.1, and the closedness of $\mathcal{C}_{d}$ is established in Proposition 4.4.
To each polynomial in $\mathcal{A}_{d}$, we associate a vector of its coefficients as an element in $\mathcal{B}(\mathcal{H})^{\langle x\rangle_{d}}.$ The topology on $\mathcal{A}_{d}$ is then the topology induced from the product WOT on $\mathcal{B}(\mathcal{H})^{{\langle x\rangle_{d}}}.$ Alternately, in this topology a net $(p_{\alpha}=\sum_{w}P_{\alpha,w}w)_{\alpha}$ in $\mathcal{A}_{d}$ converges to $p=\sum_{w}P_{w}w\in\mathcal{A}_{d}$ if for each $w\in\langle x\rangle_{d}$, the net of operators $(P_{\alpha,w})$ converges to $P_{w}$ in WOT. Note that $\mathcal{A}_{d}$ becomes a locally convex topological vector space with this topology.

### v2

@@ Abstract.
The main technical input is a canonical tuple $A$ of self-adjoint operators and, in the unitary case, a canonical tuple $U$ of unitaries, both constructed from the left-regular representation on Fock space. We prove that, up to a universal constant, the norms $\|p(A)\|$ and $\|p(U)\|$ bound the operator norm of any positive semidefinite Gram matrix $G$ representing the sos polynomial $p$. This uniform control is the key input in showing that the cone of (sums of) squares is closed in the product ultraweak topology on the coefficients. A separate approximation argument then produces a separating functional that is continuous for the weak operator topology (WOT). This two-step passage between the ultraweak and WOT topologies constitutes our separation argument and yields the required WOT closedness of the sos cone. With this in hand, the GNS construction associates to such a separating linear functional a finite-rank positive semidefinite noncommutative Hankel matrix and, on its range, produces the desired tuple $Y$.
@@ 1.1. Notation
Equip $\mathcal{A}$ with the involution : on letters, $x_{j}^{*}=x_{j},$ on a word $w=x_{i_{1}}\cdots x_{i_{n}}\in\langle x\rangle,$
@@ 1.1.1. Trigonometric polynomials
for all $p\in\mathscr{A}$. The notions of length of a word, degree of a polynomial, etc. extend naturally to $\mathscr{A}$ and we let, for positive integers $d,$
@@ 1.2. Main results
The passage to operator coefficients necessitates several novel results and constructions that we expect to be of independent interest. At a high level, the proofs of Theorem 1.1 and Theorem 1.3 still follow the now standard paradigm for establishing sum of squares (sos) representations (factorizations). Namely, the Hahn-Banach theorem produces a separating linear functional $\varphi$, and then a Gelfand-Naimark-Segal (GNS) construction based on $\varphi$ ultimately produces a tuple $Y$. Here we roughly follow the outline of [MP05].
A key construction is that of a tuple of self-adjoint operators $A$ based upon the left regular representation on Fock space; see Section 3. We then show that, up to a universal constant, for a sum of squares polynomial $p,$ the norm of $p(A)$ bounds the norm of any non-commutative psd Gram matrix $G$ that represents $p$; see Proposition 3.3. This uniform bound is the main input in Section 4 for proving that the cone $\mathcal{C}_{d}$ of sums of squares is closed in the product ultraweak topology on the coefficients. A separate approximation argument then replaces an ultraweak continuous separating functional by a WOT continuous one and hence yields closedness of $\mathcal{C}_{d}$ in the product WOT. This two-step interplay between the ultraweak and WOT topologies enables an application of the Hahn-Banach separation theorem.
On the GNS side, we introduce a new argument that exploits the WOT to associate to a separating linear functional a finite-rank psd noncommutative representing Hankel matrix and, on its range, construct the desired tuple $Y$; see Section 5. For the unitary result, Theorem 1.3, we additionally modify the construction of Section 3 to produce a canonical tuple of unitary operators from the left-regular representation and adapt the GNS procedure to obtain a unitary tuple together with a representing vector that realizes the separating functional as a vector state; see Subsection 7.2.∎
@@ 1.3. Reader’s guide
The paper is structured as follows. The convex cone of sums of squares (making an appearance in (1.3) of Theorem 1.1) is introduced and characterized in the next Section 2. In Section 3 we define creation operators $L_{i}$ on the full Fock $\mathcal{F}_{\tt{g}}^{2}$ space and their symmetrized analogs $A_{i}$. How they pertain to the sum of squares statement at hand is explored in Subsection 3.3, where evaluations at $A$ are used to extract coefficients of a polynomial. In Section 4 we introduce a suitable topology on $\mathcal{A}{d}$ and collect all the necessary topological properties needed in the sequel. With respect to this topology, the convex cone of sums of squares is closed, see Proposition 4.1. The fact that the cone is closed allows for an application of the Hahn–Banach Separation Theorem, which is then followed by an appropriate version of the GNS construction, carried out in Section 5; see Proposition 5.4. Then Theorem 1.1 is proved in Section 6 and Theorem 1.3 is proved in Section 7.
Acknowledgment
The authors thank Dr. Matthias Schötz for carefully reading an earlier version of this manuscript and for pointing out gaps in the proof of Proposition 4.1.
@@ 2. Convex Cone of (Sums of) Squares
In this section a key player in the proof of Theorem 1.1, the convex cone of sums of squares of polynomials, is introduced and studied. The main result in this section is Proposition 2.2 (see also Remark 2.4), which gives a bound on the number of sums of squares needed to write a polynomial as a sum of squares.
@@
so that $r^{*}r$ has a representation as $V_{d}^{*}GV_{d}$ for a psd matrix $G.$
@@
A polynomial $p\in\mathcal{A}_{2d}$ is in $\mathcal{C}_{d}$ if and only if there is a psd block matrix $G$ such that
$p\ =\ V_{d}^{*}GV_{d}.$ (2.2)
In fact, if $p=V_{d}^{*}GV_{d},$ then factoring $G=\sum_{j=1}^{N(d)}R_{j}R_{j}^{*}$ with $R_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ as in Lemma 2.1, setting $r_{j}=R_{j}^{*}V_{d}$ gives,
@@
We call any psd block matrix $G$ satisfying equation (2.2) a Gram representation for $p.$
@@
Given a sum of squares $p=\sum_{i=1}^{N(d)}r_{i}^{*}r_{i},$ writing $r_{j}=R_{j}^{*}V_{d}$ gives $p=\sum_{i=1}^{N(d)}V_{d}^{*}R_{i}R_{i}^{*}V_{d},$ where $R_{i}$ is the coefficient vector corresponding to the polynomial $r_{i}.$ It follows that $p=V_{d}^{*}GV_{d},$ where $G=\sum_{i=1}^{N(d)}R_{i}R_{i}^{*}.$ In particular, $G:\mathcal{H}^{N(d)}\to\mathcal{H}^{N(d)}$ is a psd linear map.
Conversely, suppose there is a psd linear map $G:\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that $p=V_{d}^{*}GV_{d}.$ By Lemma 2.1, there exist $R_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that $G=\sum_{j=1}^{N(d)}R_{j}R_{j}^{*}.$ Setting $r_{j}=R_{j}^{*}V_{d}$, one obtains $p=\sum_{j=1}^{N(d)}r_{j}^{*}r_{j}.$
By what has already been proved, if $p,q\in\mathcal{C}_{d},$ then there exist (psd) Gram representations $p=V^{*}G_{p}V$ and $q=V^{*}G_{q}V.$ Now $p+q=V_{d}^{*}(G_{p}+G_{q})V_{d}.$ Since $G_{p}+G_{q}:\mathcal{H}^{N(d)}\to\mathcal{H}^{N(d)}$ is a psd linear map, what has already been proved shows $p+q\in\mathcal{C}_{d}.$
@@
$\mathcal{C}_{d}=\left\{V_{d}^{*}GV_{d}:\quad G=[G_{v,w}]_{v,w\in\langle x\rangle_{d}}\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)},\quad G\succeq 0\right\}.$
@@ 3. Full Fock Space and Gram Matrices
This section recalls the well-known definition of the full Fock space [AP95, JMS21], the creation operators [Fra84], and introduces their symmetrized variants in (3.4) compressed to a suitable finite-dimensional subspace. In Subsection 3.3 we explore how these self-adjoint operators are used to extract the coefficients of a polynomial. The main result is Proposition 3.3 showing that the set of positive semidefinite Gram matrices of a polynomial is norm bounded.
@@ 3.2. Left creation operators
Clearly, each $L_{i}$ is an isometry. Moreover $L_{i}^{*}L_{j}=0$ if $i\neq j.$ Thus the tuple $L=(L_{1},\dots,L_{{\tt{g}}})$ is a row isometry. Let $B$
$B_{i}\ =\ L_{i}+L_{i}^{*},\quad i=1,\dots,{\tt{g}}.$
Fix a positive integer $\lcal{d}$. Let $\mathcal{F}^{2}_{{\tt{g}},\lcal{d}}$ denote the subspace of $\mathcal{F}^{2}_{{\tt{g}}}$ spanned by $\{e_{w}:w\in\langle x\rangle_{\lcal{d}}\}$ and $\iota=\iota{d}:\mathcal{F}^{2}_{{\tt{g}},\lcal{d}}\to\mathcal{F}^{2}_{{\tt{g}}}$ the inclusion. Thus, for instance, for $|w|\leq\lcal{d},$
$\iota^{*}L_{i}\iota e_{w}=\iota^{*}L_{i}e_{w}=\begin{cases}e_{x_{j}w}&\text{ if }|w|<\lcal{d}\\ 0&\text{ if }|w|=\lcal{d}.\end{cases}$ (3.2)
Similarly, if $|v|\leq\lcal{d},$ then
$\iota^{*}L_{j}^{*}\iota e_{v}=\iota^{*}L_{j}^{*}e_{v}=\begin{cases}e_{u}&\text{ if }v=x_{j}u\\ 0&\text{otherwise}.\end{cases}$ (3.3)
Let AA
$A_{j}=\iota^{*}B_{j}\iota=\iota^{*}\left(L_{j}+L_{j}^{*}\right)\iota.$ (3.4)
From equations (3.2) and (3.3), $A_{j}w=B_{j}w$ for $|w|<\lcal{d}.$ While $A$ depends on $\lcal{d}$, we suppress this dependence in the notation for readability; when $\lcal{d}$ is not clear from the context, we write $A=A^{(\lcal{d})}$.
@@
The proof proceeds by induction on word length. If $w$ is the empty word, then $A^{w}\Omega=\Omega=e_{\varnothing}.$ Assume $n\leq\lcal{d}$ and the claim holds for all words of length $\leq n-1$. Let $w=x_{i_{1}}\dots x_{i_{n}}=x_{i_{1}}\widetilde{w}$ be of length $n.$ Then
$A^{w}\Omega\ =$ $\ L_{i_{1}}A^{\tilde{w}}\Omega+L_{i_{1}}^{*}A^{\tilde{w}}\Omega,$
By the induction hypothesis,
@@
The $N(\lcal{d})\times N(\lcal{d})$ scalar matrix
$\mathbb{E}_{\lcal{d}}\ =\ \begin{bmatrix}\langle A^{w}\Omega,e_{v}\rangle\end{bmatrix}_{v,w\in\langle x\rangle_{\lcal{d}}}$
@@
Recall that we have endowed $\langle x\rangle$ with graded lexicographic order. If $\lcal{d}\geq|v|$ and $v>w$, then
@@
by Lemma 3.1. Hence, $\mathbb{E}{d}$ is upper triangular. Moreover, each diagonal entry is 11 by Lemma 3.1. Thus, $\mathbb{E}_{\lcal{d}}$ is invertible. ∎
@@ 3.3. Extraction formula for coefficients
Let $q=\sum Q_{w}w\in\mathcal{A}_{\lcal{d}}.$ For $v\in\langle x\rangle_{\lcal{d}}$ define the linear functional $\Omega_{v}:\mathcal{B}(\mathcal{F}^{2}_{{\tt{g}},\lcal{d}})\to\mathbb{C}$ by
$\Omega_{{v}}(T)=\langle T\Omega,e_{v}\rangle.$
@@
$\ =$ $\ \sum\limits_{w}\langle A^{w}\Omega,e_{v}\rangle\,Q_{w}\ =\ \sum\limits_{w}[\mathbb{E}_{\lcal{d}}]_{v,w}Q_{w},$
where $\mathbb{E}_{\lcal{d}}]_{v,w}$ is the $(v,w)$ entry of the matrix $\mathbb{E}_{\lcal{d}}.$ In short,
[…]
where $Z(q)$ and $Q$ are column vectors with $Z_{v}(q)$ and $Q_{v}$ as the $v^{\rm th}$ entry of $Z$ and $Q,$ respectively. Since, by Lemma 3.2, $\mathbb{E}_{d}$ is invertible,
[…]
We refer to $\mathbb{E}$ as the extraction matrix, and equation (3.6) as the extraction formula for the coefficients of $q.$ Note that this formula depends only upon $q(A);$ that is, the coefficients of $q$ are determined uniquely by $q(A).$
It follows from equation (3.6) that there exists a positive constant $\lambda_{\lcal{d}}$ (independent of $q$) such that
$\|Q_{w}\|\ \leq\ \lambda_{\lcal{d}}\,\|q(A)\|\quad\text{ for all $w\in\langle x\rangle_{\lcal{d}}$.}$ (3.7)
@@
$\Gamma_{p}\ =\ \{\,G\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}\ :\quad G\succeq 0,\quad V_{d}^{*}GV_{d}=p\,\}$
is norm bounded ((with respect to the operator norm on $\mathcal{B}(\mathcal{H}^{N(d)})$)). More precisely, there exists a constant $\mu_{d}$ ((depending only on $d$ and $\tt{g}$} and not on $p$)) such that, for all $G\in\Gamma_{p},$
$\|G\|\ \leq\ \mu_{d}\;\|p(A)\|,$
where the tuple $A=A^{(\lcal{d})}$ is defined in (3.4) for any $\lcal{d}\geq 2d$.
@@
Fix $p\in\mathcal{C}_{d}$ and $G\in\Gamma_{p}.$ Thus $p=V_{d}^{*}GV_{d}.$ By Proposition 2.2, there exists $Q_{j}:\mathcal{H}\to\oplus_{w\in\langle x\rangle_{d}}\mathcal{H}$ such that
@@
$\sum_{u,v\in\langle x\rangle_{d}}\|G_{u,v}\|\ \leq\ \sum_{u,v\in\langle x\rangle_{d}}\,\sum_{j=1}^{N(d)}\|Q_{j,u}Q_{j,v}^{*}\|\ \leq\ N(d)^{3}\lambda_{d}^{2}\;\|p(A)\|.$
It follows that $\|G\|\leq\mu_{d}\,\|p(A)\|$ for $\mu_{d}=N(d)^{3}\lambda_{d}^{2}.$
4. Topology on $\mathcal{A}_{\lcal{d}}$
The main purpose of this section is to define a well-behaved topology on $\mathcal{A}_{\lcal{d}}$ in which the convex cone of sums of squares is closed.
To each polynomial in $\mathcal{A}_{\lcal{d}}$ we associate the vector of its coefficients as an element in $\mathcal{B}(\mathcal{H})^{\langle x\rangle_{\lcal{d}}}.$ The topology on $\mathcal{A}_{\lcal{d}}$ is then the topology induced from the product WOT on $\mathcal{B}(\mathcal{H})^{{\langle x\rangle_{\lcal{d}}}}.$ Alternately, in this topology a net $(p_{\alpha}=\sum_{w}P_{\alpha,w}w)_{\alpha}$ in $\mathcal{A}_{\lcal{d}}$ converges to $p=\sum_{w}P_{w}w\in\mathcal{A}_{\lcal{d}}$ if and only if for each $w\in\langle x\rangle_{\lcal{d}}$, the net of operators $(P_{\alpha,w})_{\alpha}$ converges to $P_{w}$ in WOT. Note that $\mathcal{A}_{\lcal{d}}$ becomes a locally convex topological vector space with this topology. We refer to this topology as the WOT on $\mathcal{A}_{\lcal{d}}$ or the product WOT topology. It is the default topology on $\mathcal{A}_{\lcal{d}};$ that is, unless otherwise stated, it is the topology on $\mathcal{A}_{\lcal{d}}.$
Exclusive to this section, we will also use a product ultraweak topology on $\mathcal{A}_{\lcal{d}}.$ The ultraweak topology on $\mathcal{B}(\mathcal{H})$ is the weak-* topology on $\mathcal{B}(\mathcal{H})$ induced by the predual $\mathcal{T}(\mathcal{H})$, the trace class operators on $\mathcal{H}.$ It is the weakest topology such that predual elements remain continuous on $\mathcal{B}(\mathcal{H}).$ As before, to each polynomial in $\mathcal{A}_{\lcal{d}}$, we associate a vector of its coefficients as an element in $\mathcal{B}(\mathcal{H})^{\langle x\rangle{d}}.$ The ultraweak topology on $\mathcal{A}{d}$ is then the topology induced from the product ultraweak topology on $\mathcal{B}(\mathcal{H})^{{\langle x\rangle{d}}}.$ Alternately, in this topology a net $(p_{\alpha}=\sum_{w}P_{\alpha,w}w)_{\alpha}$ in $\mathcal{A}_{\lcal{d}}$ converges to $p=\sum_{w}P_{w}w\in\mathcal{A}_{\lcal{d}}$ if and only if for each $w\in\langle x\rangle_{\lcal{d}}$, the net of operators $(P_{\alpha,w})_{\alpha}$ converges to $P_{w}$ in ultraweak topology for all $w\in\langle x\rangle_{\lcal{d}}$. Note that $\mathcal{A}_{\lcal{d}}$ becomes a locally convex topological vector space with this topology.
4.1. Closedness of the cone $\mathcal{C}_{d}$
The main goal of this subsection is to establish the following result. In the rest of this section fix an integer $\lcal{d}\geq 2d$, and set $A=A^{(\lcal{d})}$.